Mrs Chen is six times as old as her son now. If she is 36 years old now, in how many years’ time will she be three times her son’s age?
1
Expert's answer
2012-09-28T08:12:02-0400
Her son is 36/6=6 years old now. In an x year she'll be 36+x and he'll be 6+x she'll be three times older, so (36+x)/(6+x)=3

36+x=18+3x

2x=18

x=9

Answer - in a 9 years.